OpenStudy (anonymous):

The sum of a number and twelve times its reciprocal is -7. find the number.

OpenStudy (anonymous):

i have it setup as x+12/x=-7

OpenStudy (anonymous):

x+12/x=-7 x^2+12=-7x x^2+7x+12=0 (x+3)(x+4)=0 x=-3, -4

OpenStudy (anonymous):

yes you r right

OpenStudy (anonymous):

we get a quadratic eq,

OpenStudy (anonymous):

ok i see i was not factoring correctly now